OpenStudy (thecalchater):

Find the value of the derivative (if it exists) of the function f(x)=7-|x| at the extrema point (0,7). A. Does not exist B. 0 C. 7 C. -7 Find any critical numbers of the function f(0)=2 sec0 + tan0,10pi<0<12pi the 0 is delta A. 67pi/6 B. 71pi/6 C.63pi/6 D. a and b

OpenStudy (anonymous):

The first one doesnt exist...normally like that for absolute value problems

OpenStudy (anonymous):

next one is D solve for 2sec10pi + tan 10pi and 2sec 12pi + tan 12pi

OpenStudy (thecalchater):

Thanks can you help with a few more ill be sure to medal and fan. @JuanitaM

OpenStudy (thecalchater):

Locate the absolute value extrema of the function f(x)=cos(pi x) on the closed interval [0,3/4] Determine whether Rolle's Theorem can be applied to the function f(x)=(x+3)(x-2)^2 on the closed interval [-3,2] if rolle's Theorem can be applied find all numbers c in the open interval (-3.2) such that f' (c)=0 Determine whether the Mean Value Theorem can be applied to the function f(x)=2 sin x + sin 2x on the closed interval [7 pi, 8 pi] If the mean value theorem can be applied find all numbers c in the open interval (7 pi, 8 pi) such that f'(c)= f(8 pi)-f(7 pi)/8 pi-7 pi